ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 46L05 Mathieu, Martin, NUI Maynooth, Ireland A simple local multiplier algebra In this joint paper with Pere Ara (Barcelona) we present a class of approximately finite dimensional C*-algebras whose local multiplier algebras are simple with real rank zero and stable rank one. Reference: P. Ara and M. Mathieu, A simple local multiplier algebra, Math. Proc. Cambridge Phil. Soc. , to appear.